OpenStudy (richyw):

This is ridiculous but I don't see how to find the between two surfaces. I'm trying to evaluate the surface integral \[\iint_{\partial S} \mathbf{F}\cdot \mathbf{n} dA\] where \(F(x,y,x)=x\hat{i}+y\hat{j}+z\hat{k}\) and S is the boundary of the region \(x^2+y^2\leq x\leq \sqrt{2-x^2-y^2}\) oriented so that the normal points out of the region. I want to do it using the divergence theroem. So basically what I am stuck on is finding the intersection of those two surfaces because from there I would have a region for my double integral in this

OpenStudy (richyw):

\[\iint_R \int^{\sqrt{2-x^2-y^2}}_{x^2+y^2}3dzdA\]

OpenStudy (richyw):

I mean I hacked away at it like this \[(x^2+y^2)^2=2-(x^2+y^2)\]\[(x^2+y^2)(x^2+y^2+1)=2\]which makes it \(x^2+y^2=1\) but honestly that's sketchy. How do I find this is a systematic test-taking no brain way?

OpenStudy (anonymous):

whitch to polar system..

OpenStudy (richyw):

how does that make the equality more obvious?

OpenStudy (anonymous):

Wow very complicated sorry I can't help.

OpenStudy (anonymous):

in polar form, your region becomes, \[r^2\le r\cos\theta\le\sqrt{2-r^2}\]

OpenStudy (anonymous):

two circles .. one centered at origin and one shifted along x-axis